OpenStudy (anonymous):

\[\frac{ 3 }{ 7 }t+21=-15\] solve for t

OpenStudy (mathmale):

First, I'd suggest you subtract 21 from both sides of this eq'n, so as to simplify the equation and to isolate (3/7) t on the left side. Next, I'd suggest you multiply both sides of the resulting equation by (7/3). What happens? Of course, there are other ways in which to solve this equation for t.

OpenStudy (jdoe0001):

\(\bf \cfrac{3}{7}t+21\implies \cfrac{3t}{7}+\cfrac{21}{1}=\cfrac{}{{\color{brown}{ \textit{what's the LCD?}}}} \)

OpenStudy (phi):

I would take math male's suggestion: \[ \frac{ 3 }{ 7 }t+21=-15 \\ \frac{ 3 }{ 7 }t+21-21=-15 -21\] can you simplify that ?